# Behavior Alignment: A New Perspective of Evaluating LLM-based Conversational Recommendation Systems

### v1

@@ 1. Introduction
Recent advancements have highlighted the potential of large language models (LLMs) across a range of tasks. Their impressive recommendation performance (sileo2022zero; hou2023large; zhang2023recommendation) and language generation capability (wei2022emergent) have attracted the attention of researchers in the conversational recommender system (CRS) community (friedman2023leveraging; liao2023proactive; zhang2023user; he2023large). For example, Google Research team (friedman2023leveraging) recently proposed an LLM-based CRS that achieved huge success on YouTube. LLMs are inherently well-suited to the needs of CRS. Unlike traditional recommendation systems, which depend on user profiles and past activities, CRS prioritizes the identification of user preferences through real-time interactions (sun2018conversational; zhang2023user; friedman2023leveraging), striving to mimic the interactions of human recommenders (hayati2020inspired; li2018towards). These characteristics require strong capabilities in language understanding and generation (li2018towards; zhou2020improving), areas where LLMs excel. Consequently, there is an increasing trend of adopting LLMs within the CRS domain.
@@ 3. Behavior Alignment: the Metric
Users tend to find the conversations more natural and engaging with human recommenders. And those conversations often lead to better recommendations compared with automatic conversational recommendation systems (hayati2020inspired). This is largely because human recommenders usually deploy various sociable and preference elicitation strategies to grab more personal information of users (hayati2020inspired). In addition, we and existing studies (liao2023proactive; zhang2023user) consistently observed a substantial disparity between human’s strategies and those employed by LLM-based CRS where latter often appears naive and passive. Therefore, human recommenders’ behaviors could become valuable references for the LLM-based CRS.
@@ 4.1. Agreement with Human Preferences
An ideal evaluation metric should yield ratings that highly correlate to human preferences (deriu2021survey). We use Cohen’s Kappa to measure such correlation. The major reason for using Cohen’s Kappa instead of other measures of agreements is it explicitly accounts for the agreement that occur by change. By calculating the expected agreement under the assumption of independence between metrics and human raters, Cohen’s kappa adjusts the observed agreement accordingly. This provides a more accurate and robust measure of the true alignment by isolating agreement that exceeds chance levels. In addition to Cohen’s Kappa, we employed the Bootstrap method to establish the 2.5% to 97.5% confidence interval.
@@ 7. Acknowledgements
8. Future Work
We believe there are at least two promising directions for future work that are interesting to explore. First, we can develop a more accurate Behavior Alignment metric that adjusts penalties according to the different conversational stages faced by the conversational system. Intuitively, in the early stages of a conversation, there are fewer sentences to condition on, so the system’s behavior should be relatively random. In this case, a lighter penalty should be applied if the system’s behavior deviates from the human reference compared to later stages. One possible approach is to include a weight for each penalty, where the weight is the inverse of the entropy for the current stage. For example, we can estimate the distribution under the conditional probability $P(B_{i}|{B_{i-1},B_{i-2},...,B_{i-t}})$ by considering the most recent $t$ behaviors. This distribution can be estimated using historical data.
Another interesting direction is to use a neural network to learn the pattern of behaviors. This sequential model would guide LLM to deploy the most appropriate behavior, conditioned on the historical behaviors of both users and the system. One possible approach is to treat each behavior as a special token and train a recurrent neural network to generate the next token, similar to how language models operate.

### v2

# Behavior Alignment: A New Perspective of Evaluating LLM-based Conversational Recommender Systems

@@ 1. Introduction
Recent advancements have highlighted the potential of large language models (LLMs) across a range of tasks. Their impressive recommendation performance (sileo2022zero; hou2023large; zhang2023recommendation) and language generation capability (wei2022emergent) have attracted the attention of researchers in the conversational recommender system (CRS) community (friedman2023leveraging; liao2023proactive; zhang2023user; he2023large). For example, Google Research team (friedman2023leveraging) recently proposed an LLM-based CRS that achieved huge success on YouTube. LLMs are inherently well-suited to the needs of CRS. Unlike traditional recommender systems, which depend on user profiles and past activities, CRS prioritizes the identification of user preferences through real-time interactions (sun2018conversational; zhang2023user; friedman2023leveraging), striving to mimic the interactions of human recommenders (hayati2020inspired; li2018towards). These characteristics require strong capabilities in language understanding and generation (li2018towards; zhou2020improving), areas where LLMs excel. Consequently, there is an increasing trend of adopting LLMs within the CRS domain.
@@ 3. Behavior Alignment: the Metric
Users tend to find the conversations more natural and engaging with human recommenders. And those conversations often lead to better recommendations compared with automatic conversational recommender systems (hayati2020inspired). This is largely because human recommenders usually deploy various sociable and preference elicitation strategies to grab more personal information of users (hayati2020inspired). In addition, we and existing studies (liao2023proactive; zhang2023user) consistently observed a substantial disparity between human’s strategies and those employed by LLM-based CRS where latter often appears naive and passive. Therefore, human recommenders’ behaviors could become valuable references for the LLM-based CRS.
@@ 4.1. Agreement with Human Preferences
An ideal evaluation metric should yield ratings that highly correlate to human preferences (deriu2021survey). We use Cohen’s Kappa to measure such correlation. In addition to Cohen’s Kappa, we employed the Bootstrap method to establish the 2.5% to 97.5% confidence interval.